# CVBench: Evaluating Cross-Video Synergies for Complex Multimodal Understanding and Reasoning

### v2

@@ 1 Introduction
The ability to reason across multiple video streams is fundamental to real-world applications, such as multi-camera surveillance systems, distributed procedural learning, and multi-view activity analysis. However, despite significant progress in single-video understanding by multimodal large language models (MLLMs) maaz2023video (33, 60, 42, 15, 21, 26, 9), their capacity for cross-video relational understanding remains largely unexplored. This capability requires reasoning about entities, events, and their interrelationships across distinct spatiotemporal contexts. The research gap stems not from the insufficient use of single-video benchmarks li2024mvbench (23, 10, 12), but from inherent architectural limitations in current evaluations that enforce isolated video processing.
@@
Current MLLMs face fundamental barriers in cross-video scenarios: 1) Inter-video context linking: models lack mechanisms to persistently integrate and retain object/event states across videos separated by time and space; 2) Entity disambiguation: overlapping or visually similar entities across videos lack unique spatiotemporal grounding, causing identification errors; 3) Temporal-causal modeling: existing models poorly capture long-range dependencies and causal links spanning video sequences.
To bridge this gap and diagnose architectural limitations, we introduce CVBench, the first diagnostic benchmark for cross-video understanding and reasoning abilities. As shown in Fig. 1, it evaluates large multimodal models through three hierarchical tasks: cross-video object association for verifying entity persistence across varying visual contexts, cross-video event association for modeling temporal sequences and causal relationships across videos, and cross-video complex reasoning for integrating visual cues and external knowledge for hierarchical spatiotemporal understanding. CVBench comprises 1,000 QA pairs sampled from five diverse domains (artistic performances, sports competitions, films and television, life records, and knowledge) to ensure generalization. This domain diversity challenges models to reason across dynamic contexts, requiring spatio-temporal integration and multi-view synthesis. Expert annotations capture real-world complexities by modeling entities, events, and inter-video relationships. This rigorous foundation enables diagnostic evaluation of cross-video reasoning capabilities.
@@
Figure 1: Illustration of CVBench task definiton. CVBench focuses on three core tasks that are crucial for real-world applications: (A) cross-video object association (identifying shared entities across different videos), (B) cross-video event association (linking temporal or causal event chains that span across video sequences), and (C) cross-video complex reasoning (integrating commonsense to understand and reason about complex interrelationships between video content). The benchmark format is multi-choice QA, and here we only show the correct answer for visualization.
@@
We evaluate over ten state-of-the-art closed-sourced and open-sourced MLLMs, including GPT-4o, Gemini series, and Qwen2.5-VL, under zero-shot and chain-of-thought prompting paradigms. Our evaluation reveals significant performance gaps in these models. Specifically, GPT-4o achieves only 60% accuracy on causal reasoning tasks, which is lower than the human-level performance of 91%. Additionally, these models struggle with inter-video context retention and difficulties in disambiguating overlapping entities, which hinders their ability to integrate information from multiple sources. These findings underscore the limitations of current MLLMs in handling complex cross-video tasks, highlighting the need for more advanced architectures to address challenges like long-term context retention and entity disambiguation.
@@
  • We introduce CVBench, a diagnostic benchmark explicitly targeting cross-video understanding through object association, event association, and complex reasoning.
  • We release a carefully curated, multi-domain multiple-choice QA dataset spanning five settings to stress-test spatiotemporal integration and multi-view synthesis.
  • We provide a systematic evaluation of leading MLLMs on CVBench, exposing performance gaps and critical challenges in cross-video reasoning, and further identifying key research directions to advance the field.
@@ 2 Related Work
2.1 Multimodal Large Language Models
The evolution of MLLMs has progressed from image-based models liu2024visual (30, 8, 64, 62, 19, 29, 46) to video-capable architectures shu2023audio (41, 4, 50, 22, 51, 1, 56, 63, 61, 14, 45, 65, 7), marked by two critical innovations: 1) Long-range context modeling that enables efficient handling excessively long temporal contexts li2024videochatflash (25, 42); 2) Spatio-temporal fusion that integrates both appearance and motion dynamics with language reasoning ren2024timechat (39, 34, 59). Despite these advances, current MLLMs remain architecturally constrained to single-video scenario, lacking mechanisms for cross-video interactions, exhibiting inadequate spatiotemporal grounding for overlapping entities, and demonstrating limited modeling of temporal event chains. These fundamental limitations restrict their applicability to multi-camera surveillance and cross-video applications.
2.2 Video Benchmarks
A comprehensive and objective benchmark is essential for evaluating MLLMs, enabling the comparison and investigation of the performance of various models. Recent advances in multimodal learning have led to the development of diverse benchmarks lei2018tvqa (20, 24, 49, 53, 18, 58, 35, 5, 31, 11, 28, 38, 44, 13, 2, 16, 40, 52, 23, 37, 55) for video understanding tasks. However, existing video evaluation frameworks exhibit critical limitations in assessing cross-video reasoning capabilities. First, general multimodal frameworks are predominantly with broad domain coverage, like Video-MMMU hu2025video (17) and Video-MME fu2024video (10)), aiming to evaluating instruction-following capabilities but lack cross-video relational tasks. Second, long-term video benchmarks contains longer videos with multiple sub-events, however are confined to single source or similar creators, such as Egoschema egoschema2023 (36), TempCompass tempcompass2024 (32), and LVBench wang2024lvbench (47). These shortcomings significantly constrain the applicability of existing benchmarks to real-world settings, such as multi-camera surveillance and cross-video procedural learning, where robust cross-domain cross-video reasoning and relational understanding are critical. To address these gaps, we introduce CVBench, the first diagnostic framework for cross-video relational reasoning. CVBench pioneers a three-tiered evaluation of spatiotemporal integration capabilities, encompassing object-level association, event-level association and complex reasoning.
@@ 3.1 Task Definition
  • Object association: Cross-video object recognition, multi-video attribute recognition, joint-video counting, and cross-video entity matching. These require consistent identification of entities across domains or viewpoints despite occlusion or deformation.
  • Event association: Cross-video anomaly detection, scene recognition, key-action recognition, event retrieval, multi-view scene understanding, and temporal reasoning. These demand fusion of spatiotemporal signatures across videos to reconstruct event dynamics.
  • Complex reasoning: Joint-video spatial navigating, video difference captioning , cross-video counterfactual reasoning, joint-video summarization, and procedural transfer. These necessitate higher-order synthesis including spatial imagination, fine-grained reasoning, causal inference and adaptive knowledge transfer.
@@ 3.2 Dataset Collection
Video curation.
We employ task-driven sampling from YouTube to construct diverse video groups (at most four videos per group). Detailed statistic are shown in Fig. 3. We ensure source diversity through the inclusion of multi-creator channels (videos with similar or same creators), cross-domain associations (videos across different domain but with shared topics or intrinsic relations), and the segmentation of long videos (sequential occurrence of event-level activities). Specifically, we measure video quality under the following four curation criteria:
  • Multi-domain coverage (daily life, sports, surveillance);
  • Intrinsic inter-video relationships (differences, similarities, co-occurrence, causality causal, temporal order, etc);
  • Duration constraints (primarily within 10 min, with videos exceeding 10 min comprising less than 5% of the dataset);
  • Visual quality and availability (at least 720p 24fps videos).
@@
TempCompass tempcompass2024 (32) 410 500 1 11.4
MSVD-QA xu2017video (53) 504 13,157 1 9.8
Video-MME fu2024video (10) 900 2,700 1 1017.9
NExT-QA xiao2021next (52) 1,000 8,564 1 39.5
AutoEval-Video chen2024autoeval (5) 327 327 1 14.6
EgoSchema egoschema2023 (36) 5,063 5,063 1 180.0
ActivityNet-QA ActivityNet2019 (57) 800 8,000 1 111.4
@@
QA generation.
@@
Qwen2.5-Omni-7B xu2025qwen25omnitechnicalreport (54) 52.4 70.9 22.7 31.0 56.4 57.7 75.0 60.8
Qwen2.5-VL-7B bai2025qwen25vltechnicalreport (3) 51.3 80.0 22.7 26.2 50.9 55.8 69.2 60.8
LLaVA-OneVision-7B li2024llama (27) 52.6 83.6 40.0 38.1 45.5 42.3 61.5 52.9
VideoLLaMA3-7B damonlpsg2025videollama3 (4) 57.6 78.2 29.3 35.7 60.0 59.6 80.8 64.7
InternVL2.5-8B chen2024internvl (6) 59.4 83.6 26.7 50.0 60.0 69.2 67.3 68.6
Phi-4-Multimodal-5B abouelenin2025phi (1) 49.7 69.1 26.7 31.0 43.6 57.7 65.4 58.8
Internvideo2.5-8B wang2022internvideo (48) 57.3 85.5 33.3 47.6 56.4 55.8 63.5 64.7
@@
Video-R1-7B videor1 (9) 49.2 74.5 22.7 28.6 47.3 48.1 76.9 52.9
Table 2: Performance of MLLMs on CVBench in cross-video complex reasoning tasks, evaluated across closed-source and open-source MLLMs. The tasks include: multi-view scene understanding (M. SU), multi-video temporal reasoning (M. TR), joint-video spatial navigation (J. SN), video difference captioning (VDC), cross-video counterfactual reasoning (C. CR), joint-video summarization (J. S), and cross-video procedural transfer (C. PT). For human evaluation, we employed five annotators and reported the average accuracy.
Quality control.
3.3 Comparison with Previous Benchmarks
4 Benchmarked Results
4.1 Experimental Settings
Models.
Our evaluation comprehensively assesses various MLLMs, including closed-source and open-source models. A prominent model in our suite is Gemini team2023gemini (43), a closed-source model known for its sophisticated multimodal processing capabilities. On the open-source side, Qwen2.5-VL videor1 (9) stands out, offering transparency with its accessible codebase for further exploration and customization. We also evaluate advanced multimodal reasoning models such as Video-R1 videor1 (9), which leverages reinforcement learning to integrate multimodal information and perform complex cross-video reasoning tasks. For ablation study, we choose Qwen2.5VL-7B as the default model.
Settings.
@@ 4.1 Experimental Settings
Qwen2.5-Omni-7B xu2025qwen25omnitechnicalreport (54) 54.4 52.9 60.9 35.0 59.5
Qwen2.5-VL-7B bai2025qwen25vltechnicalreport (3) 53.2 52.9 63.0 33.3 51.4
LLaVA-OneVision-7B li2024llama (27) 48.0 39.7 65.2 35.0 43.2
VideoLLaMA3-7B damonlpsg2025videollama3 (4) 53.6 45.6 73.9 35.0 50.0
InternVL2.5-8B chen2024internvl (6) 60.9 58.8 76.1 43.3 54.1
Phi-4-Multimodal-5B abouelenin2025phi (1) 48.8 55.9 56.5 30.0 40.5
Internvideo2.5-8B wang2022internvideo (48) 59.3 60.3 67.4 43.3 56.8
@@
Video-R1-7B videor1 (9) 54.4 54.4 63.0 30.0 40.5
Table 3: Performance of MLLMs on CVBench regarding cross-video object association, evaluated across closed-source and open-source MLLMs. Tasks include: cross-video object recognition (C. OR), multi-video attribute recognition (M. AR), joint-video counting (J. C), and cross-video entity match (C. EM).
@@
GPT-4o-mini 60.0 56.0 60.4 58.0 69.4
GPT-4o 70.8 66.7 74.5 65.9 75.5
Gemini-1.5-flash 68.7 64.3 70.5 64.8 77.6
Gemini-2.0-flash 67.0 65.5 67.8 67.0 67.3
@@
Qwen2.5-Omni-7B xu2025qwen25omnitechnicalreport (54) 58.9 53.6 61.1 58.0 63.3
Qwen2.5-VL-7B bai2025qwen25vltechnicalreport (3) 54.0 53.6 49.7 58.0 61.2
LLaVA-OneVision-7B li2024llama (27) 51.6 33.3 65.8 54.5 34.7
VideoLLaMA3-7B damonlpsg2025videollama3 (4) 60.3 56.0 61.1 63.6 59.2
InternVL2.5-8B chen2024internvl (6) 62.4 57.1 61.7 64.8 69.4
Phi-4-Multimodal-5B abouelenin2025phi (1) 51.9 44.0 54.4 59.1 44.9
Internvideo2.5-8B wang2022internvideo (48) 64.0 60.7 62.4 68.2 67.3
@@
Video-R1-7B videor1 (9) 61.6 59.5 65.8 55.7 63.3
Table 4: Performance of MLLMs on CVBench regarding cross-video event association, evaluated across closed-source and open-source MLLMs. Tasks include: cross-video anomaly detection (C. AD), cross-video scene recognition (C. SR), multi-video key-action recognition (M. KAR), and cross-video event retrieval (C. ER).

### v3

# CVBench: Benchmarking Cross-Video Synergies for Complex Multimodal Reasoning

@@
Abstract
While multimodal large language models (MLLMs) exhibit strong performance on single-video tasks (e.g., video question answering), their capability for spatiotemporal pattern reasoning across multiple videos remains a critical gap in pattern recognition research. However, this capability is essential for real-world applications, including multi-camera surveillance and cross-video procedural learning. To bridge this gap, we present CVBench, the first diagnostic benchmark designed to assess cross-video relational reasoning rigorously. CVBench comprises 1,000 question-answer pairs spanning three hierarchical tiers: cross-video object association (identifying shared entities), cross-video event association (linking temporal or causal event chains), and cross-video complex reasoning (integrating commonsense and domain knowledge). Built from five domain-diverse video clusters (e.g., sports, life records), the benchmark challenges models to analyze and integrate spatiotemporal patterns from dynamic visual streams. Extensive evaluation of 10+ leading MLLMs (including GPT-4o, Gemini-2.0-flash, Qwen2.5-VL) under zero-shot or chain-of-thought prompting paradigms. Key findings reveal stark performance gaps: even top models, such as GPT-4o, achieve only 63.5% accuracy on causal reasoning tasks, compared to the 91.3% accuracy of human performance. Crucially, our analysis reveals fundamental bottlenecks inherent in current MLLMs architectures, notably deficient inter-video context retention and poor disambiguation of overlapping entities. CVBench establishes a rigorous framework for advancing pattern recognition methodologies in multi-video scenarios, providing architectural insights for next-generation models. The data and evaluation code are available at: https://github.com/Hokhim2/CVBench.
@@ Abstract
keywords:
Multimodal large language models, Cross-video reasoning, Video question answering, Multivideo understanding, Benchmark evaluation.
journal: Pattern Recognition
[…]
[label1]organization=School of Systems Science and Engineering, Sun Yat-sen University,addressline=No. 135 Xingang West Road, city=Guangzhou, postcode=510275, state=Guangdong, country=China
[…]
[label2]organization=Tencent,addressline=Keji Zhongyi Road, Nanshan District, city=Shenzhen, postcode=518057, state=Guangdong, country=China
[…]
[label3]organization=School of Electronics and Communication Engineering, Sun Yat-sen University,addressline=No. 66 Gongchang Road, city=Shenzhen, postcode=518107, state=Guangdong, country=China
[…]
[label4]organization=School of Electronics and Information Engineering, Foshan University,addressline=No. 18 Jiangwan 1st Road, city=Foshan, postcode=528225, state=Guangdong, country=China
[…]
[label5]organization=Department of Computer Science and Engineering, Southern University of Science and Technology, addressline=1088 Xueyuan Avenue, city=Shenzhen, postcode=518055, state=Guangdong, country=China
{highlights}
First systematic benchmark for evaluating spatiotemporal relational pattern recognition across multiple videos, featuring a hierarchical structure with 1,000 expert-annotated QA pairs to diagnose capabilities in object association, event association, and complex reasoning.
Comprehensive evaluation reveals significant limitations of current models: even state-of-the-art multimodal models achieve only 63.5% accuracy on causal reasoning, far below human performance of 91.3%, exposing fundamental bottlenecks in inter-video pattern integration and entity disambiguation.
Provides actionable architectural insights for advancing pattern recognition models, including cross-video memory mechanisms and grounded spatiotemporal reasoning, offering a clear pathway toward robust multi-video understanding systems.
@@ 1 Introduction
The ability to reason across multiple video streams is fundamental to real-world applications, such as multi-camera surveillance systems, distributed procedural learning, and multi-view activity analysis Wu2025 . While multimodal large language models (MLLMs) have demonstrated significant progress in single-video understanding CAI2025111670 , their capacity for spatiotemporal pattern reasoning across videos remains a critical and underexplored challenge in pattern recognition research ZHENG2025111319 . This capability requires models to identify and analyze relational patterns across distinct spatiotemporal contexts 11008449 , such as entity correspondences, event sequences, and causal links. The research gap stems not merely from a lack of single-video benchmarks, but from the inherent limitations of existing evaluations in capturing cross-video relational dynamics, which are central to real-world pattern recognition systems.
@@
Current MLLMs face fundamental barriers in cross-video scenarios: 1) Inter-video context linking: models lack mechanisms to persistently integrate and retain object/event states cross-video separated by time and space Li10658165 , 11093290 ; 2) Entity disambiguation: overlapping or visually similar entities cross-video lack unique spatiotemporal grounding, causing identification errors; 3) Temporal-causal modeling: existing architectures poorly capture long-range dependencies and causal links spanning asynchronous video sequences 11095110 .
@@
Figure 1: Illustration of CVBench task definition. CVBench focuses on three core tasks that are crucial for real-world applications: a) cross-video object association (identifying shared entities across different videos), b) cross-video event association (linking temporal or causal event chains that span across video sequences), and c) cross-video complex reasoning (integrating commonsense to understand and reason about complex interrelationships between video content). The benchmark format is multi-choice QA, and here we only show the correct answer for visualization.
To address these limitations and provide a diagnostic framework for the community, we introduce CVBench, the first comprehensive benchmark designed specifically for evaluating cross-video relational pattern recognition. As illustrated in Fig. 1, CVBench structures the evaluation around three hierarchically organized tasks:cross-video object association for verifying entity persistence across varying visual contexts, cross-video event association for modeling temporal sequences and causal relationships cross-video, and cross-video complex reasoning for integrating visual cues and external knowledge for hierarchical spatiotemporal understanding. CVBench comprises 1,000 QA pairs sampled from five diverse domains (artistic performances, sports competitions, films and television, life records, and knowledge) to ensure generalization. This domain diversity challenges models to reason across dynamic contexts, requiring spatio-temporal integration and multi-view synthesis. Expert annotations capture real-world complexities by modeling entities, events, and inter-video relationships. This rigorous foundation enables diagnostic evaluation of cross-video reasoning capabilities.
@@
We evaluate over ten state-of-the-art closed-sourced and open-sourced MLLMs, including GPT-4o, Gemini series, and Qwen2.5-VL, under zero-shot and chain-of-thought prompting paradigms. Our evaluation reveals significant performance gaps in these models. Specifically, GPT-4o achieves only 60.2% accuracy on causal reasoning tasks, which is lower than the human-level performance of 91.3%. Additionally, these models struggle with inter-video context retention and difficulties in disambiguating overlapping entities, which hinders their ability to integrate information from multiple sources. These findings underscore the limitations of current MLLMs in handling complex cross-video tasks, highlighting the need for more advanced architectures to address challenges like long-term context retention and entity disambiguation.
@@
1) We introduce CVBench, the first diagnostic benchmark explicitly targeting cross-video understanding through object association, event association, and complex reasoning.
2) We release a carefully curated multi-domain dataset with comprehensive annotations, enabling rigorous stress-testing of spatiotemporal integration and multi-view synthesis capabilities.
3) We provide a systematic analysis of state-of-the-art MLLMs on CVBench, exposing critical performance gaps and identifying key research directions for advancing cross-video pattern reasoning in the field.
@@ 2 Related work
Multimodal large language models
The evolution of MLLMs has progressed from image-based Zhang10547418 to video-capable architectures Fu11093290 , ZHANG2026111925 , marked by two main improvements: 1) Long-range context modeling that enables efficient handling excessively long temporal contexts Yue11093104 , Song10657734 ; 2) Spatio-temporal fusion that integrates both appearance and motion dynamics with language reasoning Ren10656135 , Zhang11094660 , Liu11094049 . Despite these advances, current MLLMs remain architecturally constrained to single-video scenario, lacking mechanisms for cross-video interactions, exhibiting inadequate spatiotemporal grounding for overlapping entities, and demonstrating limited modeling of temporal event chains. These fundamental limitations restrict their applicability to multi-camera surveillance and cross-video applications Yang11037495 .
Video understanding benchmarks
A comprehensive and objective benchmark is essential for evaluating MLLMs, enabling the comparison and investigation of the performance of various models SHAO2026112724 . Recent advances in multimodal learning have led to the development of several benchmarks for video understanding tasks ZHONG2025111035 . However, existing video evaluation frameworks exhibit critical limitations in assessing cross-video reasoning capabilities. First, general multimodal frameworks are predominantly with broad domain coverage, such as Video-MMMU and Video-MME Fu11093290 ), to evaluate instruction-following capabilities but lacking cross-video relational tasks. Second, long-term video benchmarks contain longer videos with multiple sub-events, however, are confined to single source or similar creators, such as VideoTree Wang11094563 , Video-Bench Ning11284911 , and Chapter-Llama Ventura11092338 . These deficiencies fundamentally limit their applicability to real-world scenarios in multi-camera surveillance and cross-video procedural learning.
To address these gaps, we introduce CVBench, the first diagnostic framework for cross-video relational reasoning across different sources. CVBench pioneers a three-tiered evaluation of spatiotemporal integration capabilities, encompassing object-level association, event-level association and complex reasoning.
@@ 3.1 Task definition
1) Object association: Cross-video object recognition, multi-video attribute recognition, joint-video counting, and cross-video entity matching. These require consistent identification of entities across domains or viewpoints despite occlusion or deformation.
2) Event association: Cross-video anomaly detection, scene recognition, key-action recognition, event retrieval, multi-view scene understanding, and temporal reasoning. These demand fusion of spatiotemporal signatures across videos to reconstruct event dynamics.
3) Complex reasoning: Joint-video spatial navigating, video difference captioning, cross-video counterfactual reasoning, joint-video summarization, and procedural transfer. These necessitate higher-order synthesis including spatial imagination, fine-grained reasoning, causal inference and adaptive knowledge transfer.
@@ 3.2 Dataset collection
3.2.1 Video curation
We employ task-driven sampling from YouTube to construct diverse video groups (at most four videos per group). Detailed statistics are shown in Fig. 2. We ensure source diversity through the inclusion of multi-creator channels (videos without similar or same creators), cross-domain associations (videos across different domain but with shared topics or intrinsic relations), and the segmentation of long videos (sequential occurrence of event-level activities). Specifically, we measure video quality under the following four curation criteria: a) Multi-domain coverage, including daily life, sports, surveillance and other commen scenrios; b) Intrinsic inter-video relationships, such as differences, similarities, co-occurrence, causality causal, temporal order; c) Duration constraints, where videos are primarily within 10 minutes in length, and those exceeding 10 minutes account for less than 5% of the dataset; d) Visual quality and availability, requiring videos to have a resolution of at least 720p and a frame rate of 24 fps.
3.2.2 QA generation
3.2.3 Quality control
3.3 Comparison with previous benchmarks
@@
TempCompass tempcompass2024 410 500 1 11.4
MSVD-QA XuVideo 504 13,157 1 9.8
Video-MME Fu11093290 900 2,700 1 1017.9
NExT-QA Xiao9577425 1,000 8,564 1 39.5
AutoEval-Video chen2024Auto 327 327 1 14.6
EgoSchema egoschema2023 5,063 5,063 1 180.0
ActivityNet-QA zhou2019Activity 800 8,000 1 111.4
@@
4 Benchmarked results
4.1 Experimental settings
4.1.1 Models
Our evaluation comprehensively assesses various MLLMs, including closed-source and open-source models. A prominent model in our suite is Gemini team2023gemini , a closed-source model known for its sophisticated multimodal processing capabilities. On the open-source side, Qwen2.5-VL videor1 stands out, offering transparency with its accessible codebase for further exploration and customization. We also evaluate advanced multimodal reasoning models such as Video-R1 videor1 , which leverages reinforcement learning to integrate multimodal information and perform complex cross-video reasoning tasks. For ablation study, we choose Qwen2.5VL-7B as the default model.
4.1.2 Settings
Table 2: Performance of MLLMs on CVBench in cross-video complex reasoning tasks, evaluated across closed-source and open-source MLLMs. The tasks include: multi-view scene understanding (M. SU), multi-video temporal reasoning (M. TR), joint-video spatial navigation (J. SN), video difference captioning (VDC), cross-video causal reasoning (C. CR), joint-video summarization (J. S), and cross-video procedural transfer (C. PT).
@@ 4.1.2 Settings
Qwen2.5-Omni-7B xu2025qwen25omnitechnicalreport 52.4 70.9 22.7 31.0 56.4 57.7 75.0 60.8
Qwen2.5-VL-7B bai2025qwen25vltechnicalreport 51.3 80.0 22.7 26.2 50.9 55.8 69.2 60.8
LLaVA-OneVision-7B Li2024LLamavid 52.6 83.6 40.0 38.1 45.5 42.3 61.5 52.9
VideoLLaMA3-7B cheng2024videollama 57.6 78.2 29.3 35.7 60.0 59.6 80.8 64.7
InternVL2.5-8B Chen10656429 59.4 83.6 26.7 50.0 60.0 69.2 67.3 68.6
Phi-4-Multimodal-5B abouelenin2025phi 49.7 69.1 26.7 31.0 43.6 57.7 65.4 58.8
Internvideo2.5-8B wang2022internvideo 57.3 85.5 33.3 47.6 56.4 55.8 63.5 64.7
@@
Video-R1-7B videor1 49.2 74.5 22.7 28.6 47.3 48.1 76.9 52.9
Table 3: Performance of MLLMs on CVBench regarding cross-video object association, evaluated across closed-source and open-source MLLMs. Tasks include: cross-video object recognition (C.OR), multi-video attribute recognition (M.AR), joint-video counting (J.C), and cross-video entity match (C.EM). For human evaluation, we employed five annotators and reported the average accuracy.
@@
Qwen2.5-Omni-7B xu2025qwen25omnitechnicalreport 54.4 52.9 60.9 35.0 59.5
Qwen2.5-VL-7B bai2025qwen25vltechnicalreport 53.2 52.9 63.0 33.3 51.4
LLaVA-OneVision-7B Li2024LLamavid 48.0 39.7 65.2 35.0 43.2
VideoLLaMA3-7B cheng2024videollama 53.6 45.6 73.9 35.0 50.0
InternVL2.5-8B Chen10656429 60.9 58.8 76.1 43.3 54.1
Phi-4-Multimodal-5B abouelenin2025phi 48.8 55.9 56.5 30.0 40.5
Internvideo2.5-8B wang2022internvideo 59.3 60.3 67.4 43.3 56.8
@@
Video-R1-7B videor1 54.4 54.4 63.0 30.0 40.5
Table 4: Performance of MLLMs on CVBench regarding cross-video event association, evaluated across closed-source and open-source MLLMs. Tasks include: cross-video anomaly detection (C.AD), cross-video scene recognition (C.SR), multi-video key-action recognition (M.KAR), and cross-video event retrieval (C.ER). For human evaluation, we employed five annotators and reported the average accuracy.
@@
GPT-4o-mini 60.0 56.0 13.4 58.0 69.4
GPT-4o 70.8 66.7 14.8 65.9 75.5
Gemini-1.5-flash 68.7 64.3 14.8 64.8 77.6
Gemini-2.0-flash 67.0 65.5 14.8 67.0 67.3
@@
Qwen2.5-Omni-7B xu2025qwen25omnitechnicalreport 58.9 53.6 14.1 58.0 63.3
Qwen2.5-VL-7B bai2025qwen25vltechnicalreport 54.0 53.6 12.1 58.0 61.2
LLaVA-OneVision-7B Li2024LLamavid 51.6 33.3 12.1 54.5 34.7
VideoLLaMA3-7B cheng2024videollama 60.3 56.0 13.4 63.6 59.2
InternVL2.5-8B Chen10656429 62.4 57.1 14.8 64.8 69.4
Phi-4-Multimodal-5B abouelenin2025phi 51.9 44.0 9.4 59.1 44.9
Internvideo2.5-8B wang2022internvideo 64.0 60.7 14.1 68.2 67.3
@@
Video-R1-7B videor1 61.6 59.5 65.8 55.7 63.3